On January 1, 2021, Drenz Co. acquired 30,000 ordinarily shares out of the 100,000 outstanding ordinary shares of Josiah Inc. for ₱5,000,000. Josiah’s assets and liabilities approximate their fair values except for inventories with carrying amount of ₱800,000 and fair value of ₱900,000 and machinery with carrying amount of ₱2,500,000 and fair value of ₱2,200,000. The remaining useful life of the machinery is 5 years. Josiah’s net asset has a book value of ₱10,000,000.


On December 31,2021. Josiah reported net income of ₱2,000,000 and declared and paid dividends of ₱800,000. On December 31,2022. Josiah reported net income of ₱4,500,000 and declared and paid dividends of ₱1,600,000.


1. How much is the implied goodwill from acquisition?

2. How much is the net share in the profit or loss of the associate (investment income) in 2021?

3. How much is the carrying amount of the investment as of December 31,2021?

4. How much is the net share in the profit or loss of the associate (investment income) in 2022?


1
Expert's answer
2021-06-14T14:02:03-0400

Solution:

1.). Implied goodwill:

First, determine the percentage acquired by Drenz Co. in Josiah’s ordinary shares:

Interest acquired = "\frac{30,000}{100,000}\times 100\% = 30\%"

First derive carrying goodwill from the acquisition:

Goodwill = purchase price less net book value of assets less fair value adjustments

Purchase price                                                                                                    = 5,000,000

Book value of interest acquired = "10,000,000\times 30\%"                                 = 3,000,000

Excess cost over book value                                                                               = 2,000,000

Less: Share in undervaluation of inventory (900,000 – 800,000)"\times 30\%"      =    (30,000)

Add: Share in overvaluation of machinery (2,500,000 – 2,200,000) "\times 30\%"  =      90,000

Goodwill                                                                                                            = 2,060,000

Goodwill = 2,060,000

Derive implied goodwill:

Book value of assets inclusive of goodwill = 3,000,000 + 2,060,00 = 5,060,000

Fair value of assets = 2,200,000

Thus, the fair value of the reporting unit is less than its book value, therefore, an impairment has occurred.

 

The second step involves comparing the fair value of the assets with the book value net of goodwill and the difference is the implied goodwill amount.

Fair value of assets = 2,200,000

Book value of assets net of goodwill = 5,060,000 – 2,060,000 = 3,000,000

Implied goodwill = 3,000,000 – 2,200,000 = 800,000

Implied goodwill from acquisition = ₱800,000


2.). Net share in the profit or loss of the associate (the investment income) in 2021:

Net income reported in Dec 31, 2021 = 2,000,000

Share of profit = (2,000,000 "\times 30\%" )                                        = 600,000

Add: Overvaluation of asset adjusted = (30,000"\times 30\%" )          = 90,000

Less: Undervaluation of inventory adjusted (100,000"\times 30\%" ) = (30,000)

The net share of profit                                                                    = 660,000

The net share in the profit or loss of the associate (the investment income) in 2021 = ₱660,000


3.). Carrying amount of the investment as of December 31, 2021:

Beginning carrying amount (Cost)                                                                          = 5,000,000

Add: Share of profit (2,000,000"\times 30\%" )                                                                  = 600,000

Add: Overvaluation of asset adjusted (300,000"\times 30\%" )                                         = 90,000

Less: Cash dividends received (800,000"\times 30\%" )                                                    =  (240,000)

Less: Undervaluation of inventory adjusted (100,000"\times 30\%" )                               =   (30,000)

Ending carrying amount of the investment                                                            = 5,420,000

 

The carrying amount of the investment as of December 31, 2021, = ₱5,420,000


4.). Net share in the profit or loss of the associate (the investment income) in 2022:

Net income reported in Dec 31, 2022 = 4,500,000

Share of profit = (4,500,000"\times 30\%" )                                        = 1,350,000

The net share in the profit or loss of the associate (the investment income) in 2022 = ₱1,350,000